# DialSim: A Real-Time Simulator for Evaluating Long-Term Multi-Party Dialogue Understanding of Conversation Systems

### v5

@@ 3.2 Statistics
Table 1 shows the statistics of LongDialQA.
@@ 4 DialSim
Building on LongDialQA, DialSim features a conversation system taking on the role of a main character in a dialogue (i.e., Ross, Sheldon, and Michael). Throughout the simulation, the system is asked questions by other characters that must be answered accurately within a time limit (§ 4.1).
Input: $\mathcal{D}=\{\mathcal{S}_{i}\}_{i=1}^{N}$, Time interval $t$, System
@@
[…]
3 $T$leftarrow 0 // TotalQuestions;
4 $\mathcal{M}_{1,0}\leftarrow\phi$;
5
6for $n\leftarrow 1$ to $N$ do
@@ 11             $c\leftarrow RandCharInThreeTurns(u_{n,m})$;
12             $q_{n,m,c},a_{true}\leftarrow RandomQnA(n,m,c)$;
13             $T$leftarrow $T$ + 1;
@@
15                   $\mathcal{M}_{n,k}\leftarrow UpdateMemory(\mathcal{M}_{n,k-1},$ $u_{n,k},d_{n},t)$;
@@
17                         $a_{n,m}\leftarrow SystemAns(\mathcal{M}_{n,m},$ $q_{n,m,c},d_{n},t)$;
@@ 22                  $\mathcal{M}_{n+1,0}\leftarrow\mathcal{M}_{n,k}$;
4.1 Simulator
Algorithm 1 outlines the simulation process of DialSim, designed to emulate a real-time conversation. In this simulator, each participant’s utterance (including the system’s) occurs at a predefined time interval (same as time limit), and the system should update its memory within this interval. [^5] The memory can be incrementally updated in various ways (e.g., by storing each utterance separately or by summarizing the session up to the current utterance). A detailed discussion of these methods is provided in § 5.2. If updating the memory is not completed within the interval, the simulator will move on to the next utterance (Line 13). During the simulation, other characters ask questions (selected from LongDialQA) to the system (Line 8-10), except in sessions where the system is the only one talking (Line 5-6). The timing to ask a question is chosen randomly within the session (Line 8), and the speaker who asks the question is also chosen randomly. However, to make the simulation realistic, it is crucial to ensure that the chosen speaker is still present and hasn’t left the session. We achieved this by randomly choosing from characters who were present within three turns of the system’s last utterance (Line 9). Then, a question is randomly selected and asked in the style of the corresponding speaker (Line 10). The system then must respond to the question using its memory, all within the time limit (Line 15). The prompt for the response is created by combining the question with the dialogue history stored in the memory. If the response is not completed within the time limit, it will be considered a failure, and the simulator will move on to the next utterance. The prompt we used is provided in Appendix F.
@@ 5 Experiments
Type Model Base LLM RAG-based BM25 OpenAI Embedding Utterance Session Entire Session Sum. Utterance Session Entire Session Sum. API GPT-4o-mini 38.53 (0.89) 32.65 (2.65) 49.04 (1.67) 40.27 (1.36) 40.10 (0.75) 44.36 (2.36) 42.53 (1.26) GPT-3.5 31.82 (1.31) 25.58 (1.78) 39.70 (1.86) 32.09 (0.84) 32.06 (1.60) 36.84 (1.77) 36.69 (1.25) Gemini 1.0 pro 2.96 (0.31) 28.77 (1.83) 25.07 (2.40) 35.27 (1.80) 34.22 (0.49) 31.83 (0.41) 35.75 (2.93) Open Tülu 2-70B 0.37 (0.15) 20.94 (0.75) 20.27 (0.99) 19.75 (0.08) 31.76 (1.84) 10.15 (0.55) 18.87 (0.30) Tülu 2-7B 0.84 (0.15) 12.68 (0.24) 19.58 (1.04) 26.84 (0.85) 14.08 (0.89) 17.39 (1.37) 25.21 (1.28) Llama3.1-70B 0.60 (0.06) 31.08 (1.21) 0.55 (0.12) 16.26 (5.05) 39.00 (0.30) 2.26 (0.42) 20.14 (0.22) Llama3.1-8B 28.82 (1.94) 27.12 (0.95) 34.14 (0.85) 30.91 (0.63) 29.76 (1.31) 33.25 (0.57) 24.48 (0.60) Mixtral-8x7B 1.88 (0.26) 16.84 (0.95) 26.23 (0.90) 17.11 (1.94) 17.94 (1.32) 26.78 (1.04) 15.40 (1.39) Mistral-7B 2.82 (0.46) 24.22 (2.04) 33.07 (1.01) 29.29 (1.76) 28.30 (1.93) 29.15 (1.67) 25.41 (1.53) Gemma-7B 16.60 (0.84) 22.11 (1.73) 24.30 (2.04) 18.33 (1.37) 26.42 (2.48) 22.54 (0.78) 18.80 (0.64) Gemma-2B 0.68 (0.20) 24.06 (2.03) 24.22 (1.34) 25.79 (1.00) 25.31 (1.55) 24.48 (1.62) 25.78 (1.12) Random Guessing 20.00 20.00 20.00 20.00 20.00 20.00 20.00
$\dagger$: Both GPT-4o-mini and Llama3.1 support up to 128k tokens, but we limited them to 8k tokens due to high costs and GPU VRAM limits, respectively.
Table 2: The performance of the agents on Friends dialogue in DialSim (time limit = 6 seconds). We conducted experiments three times and reported the accuracy and standard deviations. Bold indicates the highest performance for each retrieval method.
@@ 5.1 Experimental Setting
We provide the option to choose between multiple-choice and open-ended question formats. In our experiments, we used the multiple-choice format to efficiently and accurately assess the system’s dialogue understanding capabilities. Table 1 presents the statistics for LongDialQA, highlighting a significant disparity between the number of answerable and unanswerable questions. To ensure a balanced distribution of correct answers during the simulation, 20% of the questions were intentionally designed to be unanswerable, with each question providing five possible choices.
DialSim operates in real-time, requiring precise control of the experimental environment. Therefore, we conducted all experiments using the same hardware: NVIDIA RTX A6000 GPUs and an AMD EPYC 7702 64-Core Processor. The time limit used in the experiment was set to 6 seconds, based on the average time interval between utterances in the TV shows. Note that the time limit can be set to any value (even infinity) that meets one’s service requirement. We provide extensive discussions on the time limit feature of DialSim, including the test environment control and internet speed in Appendix G, along with details about question formats.
5.2 Baselines
@@
We experimented with representative LLM-based conversation systems capable of handling long-term dialogue, focusing on two memory management approaches. The first method, namely Base LLM, simply prepends as many of the most recent utterances as the model’s context length allows. The second method, namely RAG-based, employs a retriever to search for relevant dialogue history from the system’s memory (external storage) and includes it in the prompt (Lewis et al., 2020). This method can be broken down into three ways for storing dialogue history: each speaker’s utterance individually, the entire session, and a summarized version of each session (denoted as Utterance, Session Entire, and Session Sum. in Table 2). The retrieval from the memory was performed using BM25 (Robertson et al., 2009) and cosine similarity with the OpenAI embeddings (OpenAI, 2024c).
For the LLMs, we used both API-based models (i.e., Gemini-1.0 Pro, 1.5 Pro (Team et al., 2023; Reid et al., 2024), Claude 3 Opus (Anthropic, 2024), GPT-3.5, 4o, 4o-mini (OpenAI, 2023b, 2024b, 2024a)) and open-source models (i.e., Tülu 2-7B, 70B (Ivison et al., 2023), Llama3.1-8B, 70B (Meta, 2024), Mistral-7B, 8x7B (Jiang et al., 2023, 2024), and Gemma-2B, 7B (Team et al., 2024)). [^6] Gemini-1.5 Pro, Claude 3 Opus, and GPT-4o were evaluated only in the BM25-Session Entire and oracle setting to measure their performance upper bound due to their high prices. The experimental results can be found in Appendix K. To emulate conversational settings, we used chat templates for instruction-tuned models or directly used chat models.
@@ 5.3 Results
Overall Performance   Table 2 shows that API-based models outperformed open-source models due to their superior inference capabilities and faster response times in our setting. However, the performances of all baselines were below 50%, suggesting that current LLMs have limitations in their ability to serve as conversation systems for long-term multi-party dialogues. The experimental results for Friends, The Big Bang Theory, and The Office exhibited similar trends. The detailed results are described in Appendix H.
Model Base LLM BM25 OpenAI Embedding UT SE SS UT SE SS GPT-4o-mini 38.91 34.44 49.21 42.23 38.91 43.64 42.40 GPT-3.5 31.81 26.91 39.45 32.77 32.41 35.78 35.98 Gemini 1.0 pro 28.36 28.10 39.90 34.11 34.26 30.93 33.96 Llama3.1-70B 36.36 31.84 43.17 43.81 39.85 43.17 48.49 Llama3.1-8B 28.78 29.89 34.70 33.93 31.63 32.91 35.59 Mixtral-8x7B 42.19 31.84 46.47 32.31 35.51 41.24 34.18 Mistral-7B 32.93 28.20 35.09 30.16 30.12 31.00 30.10 $\dagger$: Limited the maximum context length to 8k tokens.
Table 3: The performance of the top-performing agents on Friends in DialSim without time limit. UT, SE, and SS denote Utterance, Session Entire, and Session Summary, respectively. Bold indicates the highest performance for each retrieval method, and underlined values indicate the best-performing retrieval method for each model. The full experimental results are in Appendix I.
For real-time interactions, selecting a model size that balances inference speed and reasoning ability is crucial. As shown in Table 2, under time constraints, differences in performance between model sizes often diminish, with smaller models sometimes outperforming larger ones due to faster inference. In contrast, as detailed in Table 3, larger models generally excel when no time limits are imposed, demonstrating superior reasoning capabilities. Interestingly, larger open-source models achieve inference performance comparable to API-based models, highlighting the trade-off between speed and accuracy. Therefore, selecting a model size that achieves a balanced trade-off is critical. Additional performance comparisons under varying time constraints are provided in Appendix J.
Storing the entire session consistently outperforms other history storing methods, as shown in Table 3. This is because individual utterances lack adequate context, and crucial information may be lost during summarization. However, Llama3.1 models achieved the best performance when using Session Sum. as a history saving method, owing to their strong summarization capabilities. Additionally, contrary to our expectations, Mixtral’s Base LLM (i.e., without history retrieval) outperforms some retrieval-based models in settings with unlimited time. This is due to Mixtral’s context length of 32k tokens, which is long enough to accommodate half a season of the script, allowing it to utilize a longer dialogue history than some of the other baselines. However, in a setting with a time limit, Mixtral’s performance significantly drops due to its long inference time. Therefore, for a conversation system to converse in real-time, it is necessary to select a reasonably appropriate length of dialogue history.
Advanced techniques for storing and retrieving history are essential to engage in long-term multi-party dialogues. We conducted experiments under the oracle setting, where systems were given evidence sessions along with their dates (see Figure 2). Under these conditions, Llama3.1-70B achieved a top performance of 69.86% in an unlimited time scenario, outperforming the best RAG-based method by 21.37%. This significant performance gap highlights the importance of effective memory management techniques. Detailed experimental results are provided in Appendix K.
@@
TKG-based questions present a greater challenge than fan quiz-based ones, with two-hop questions being particularly difficult. To assess the difficulty levels across different question types, we conducted an error analysis on GPT-4o-mini, based on BM25-Session Entire, which showed the highest performance. The results showed that fan quiz-based questions had an accuracy of 58.80%, while TKG-based questions scored lower at 46.40%, highlighting the greater difficulty of TKG-based questions. Breaking down TKG-based questions further, one-hop questions had a performance of 66.67%, whereas two-hop questions had a performance of 13.53%, underscoring the challenge of two-hop questions. Furthermore, even in the oracle setting, while the performance of one-hop questions increased to 84.05%, two-hop questions remained at 28.45%. This indicates that two-hop questions are challenging not only in terms of history retrieval but also in reasoning across the given sessions.
Model Original Swapping Names New Names GPT-4o-mini 49.21 38.40 ($\downarrow$10.81) 46.57 ($\downarrow$2.64) Llama3.1-70B 48.49 39.00 ($\downarrow$9.49) 44.27 ($\downarrow$4.22) Llama3.1-8B 35.59 31.59 ($\downarrow$4.00) 32.23 ($\downarrow$3.36) Mixtral-8x7B 46.47 37.72 ($\downarrow$8.75) 39.98 ($\downarrow$6.49) Mistral-7B 35.09 30.65 ($\downarrow$4.44) 34.59 ($\downarrow$0.50)
Table 4: The performance of the top-performing open-source agents on the adversarial test (without time limit). Numbers in parentheses represent the performance drop compared to the original test. The full experimental results are provided in Appendix L.
The adversarial versions of LongDialQA are effective for the rigorous dialogue understanding evaluation of conversation systems. We conducted further experiments using DialSim on the adversarial versions of LongDialQA. The results shown in Table 4 indicated a non-trivial drop in overall performance compared to the original setup. This decline can be attributed to the systems relying not only on the dialogue history but also likely on their pre-trained knowledge when answering questions. Additionally, the performance decrease was more pronounced when names were swapped compared to when new names were assigned. This suggests that new names represent new information, while swapped names in the dialogue history conflicted with the pre-trained knowledge, leading to reduced reasoning performance. Detailed experimental results are provided in Appendix L.

### v6

@@ 3.2 Statistics
Table 2 presents the statistics of LongDialQA, highlighting a significant disparity between the number of answerable and unanswerable questions. When conducting experiments using DialSim in the form of multiple-choice questions, to ensure a balanced distribution of correct answers during the simulation, 20% of the questions were intentionally designed to be unanswerable, with each question providing five possible choices.
@@ 4 DialSim
Building on LongDialQA, DialSim features an agent taking on the role of a main character in a dialogue. Throughout the simulation, the agent is asked questions by other characters that must be answered accurately.
Algorithm 1 outlines the simulation process of DialSim, designed to emulate a conversation. In this simulator, each participant’s utterance (including the agent’s) occurs, and the agent should update its memory. [^5] The memory can be incrementally updated in various ways (e.g., by storing each utterance separately or by summarizing the session up to the current utterance). A detailed discussion of these methods is provided in § 5.2. During the simulation, other characters ask questions (selected from LongDialQA) to the agent (Line 8-10), except in sessions where the agent is the only one talking (Line 5-6). The timing to ask a question is chosen randomly within the session (Line 8), and the speaker who asks the question is also chosen randomly. However, to make the simulation realistic, it is crucial to ensure that the chosen speaker is still present and hasn’t left the session. We achieved this by randomly choosing from characters who were present within three turns of the agent’s last utterance (Line 9). Then, a question is randomly selected and asked in the style of the corresponding speaker (Line 10). The agent then must respond to the question using its memory (Line 15). The prompt for the response is created by combining the question with the dialogue history stored in the memory. The prompt we used is provided in Appendix F.
Input: $\mathcal{D}=\{\mathcal{S}_{i}\}_{i=1}^{N}$, Agent
@@
$C\leftarrow 0$ // CorrectAnswers
$T\leftarrow 0$ // TotalQuestions
2 $\mathcal{M}_{1,0}\leftarrow\varnothing$;
3
4for $n\leftarrow 1$ to NN do
@@ 9    $c\leftarrow RandCharInThreeTurns(u_{n,m})$;
10    $(q_{n,m,c},a_{true})\leftarrow RandomQnA(n,m,c)$;
11    $T\leftarrow T+1$;
@@
13       $\mathcal{M}_{n,k}\leftarrow UpdateMemory(\mathcal{M}_{n,k-1},u_{n,k},d_{n})$;
@@
15          $a_{n,m}\leftarrow AgentAns(\mathcal{M}_{n,m},q_{n,m,c},d_{n})$;
@@ 5.1 Experimental Setting
We provide the option to choose between multiple-choice and open-ended question formats. In our experiments, we used the multiple-choice format to efficiently and accurately assess the agent’s dialogue understanding capabilities. Details about the question formats and the open-ended questions can be found in Appendix G. The temperature for the LLMs was set to 0.2, and the top-p value was set to 0.1. All experiments were conducted using NVIDIA RTX A6000 GPUs and an AMD EPYC 7702 64-Core Processor.
5.2 Conversational Agents
@@
We experimented with LLM-based conversational agents capable of handling long-term dialogue, focusing on two memory management approaches. The first method, namely Base LLM, simply prepends as many of the most recent utterances as the model’s context length allows. The second method, namely RAG-based, employs a retriever to search for relevant dialogue history from the agent’s memory (external storage) and includes it in the prompt (Lewis et al., 2020). This method can be broken down into three ways for storing dialogue history: each speaker’s utterance individually, the entire session, and a summarized version of each session (denoted as Utterance, Session Entire, and Session Sum. in Table 3). The retrieval from the memory was performed using BM25 (Robertson et al., 2009) and cosine similarity with the OpenAI embeddings (OpenAI, 2024b).
For the LLMs, we used both API-based models (i.e., Gemini-2.5 Flash (Google, 2025a), Gemini-2.0 Flash (Google, 2025b), GPT-4o-mini (OpenAI, 2024a), and GPT-4.1-nano (OpenAI, 2025)) and open-source models (i.e., Llama 3.1-8B (Meta, 2024a), Llama 3.3-70B (Meta, 2024b), Mistral-7B (Jiang et al., 2023), Mixtral-8x7B (Jiang et al., 2024), Qwen 3-32B, Qwen 3-8B (Yang et al., 2025), Qwen 2.5-14B, Qwen 2.5-7B (Yang et al., 2024), Phi 4-14B, Phi 4 mini-3.8B (Abdin et al., 2024)), OLMo 2-7B (OLMo et al., 2024), OLMoE-1B-7B (Muennighoff et al., 2024), and Tülu 3-8B (Lambert et al., 2024)). To emulate conversational settings for the open-source models, we constructed chat-style prompts by applying templates using the Hugging Face apply_chat_template method. [^6] https://huggingface.co/docs/transformers/en/chat_templating
Table 3: The performance of the agents on Friends dialogue in DialSim. We conducted experiments three times and reported the accuracy and standard deviations. Bold indicates the best performance per retrieval method. Underline indicates the highest score for each model.
Type Model Base LLM RAG-based
BM25 OpenAI Embedding
Utterance Session Entire Session Sum. Utterance Session Entire Session Sum.
API GPT-4o-mini 48.11 (1.26) 35.46 (0.91) 52.11 (0.52) 44.15 (2.35) 41.29 (0.30) 45.47 (2.08) 42.78 (1.03)
GPT-4.1-nano 26.01 (0.42) 25.54 (1.13) 31.03 (2.36) 25.42 (1.48) 29.50 (2.50) 27.63 (0.97) 26.44 (0.58)
Gemini 2.5 Flash 53.94 (1.26) 30.86 (0.97) 52.92 (1.66) 24.27 (2.01) 37.16 (0.45) 42.40 (1.15) 23.75 (1.57)
Gemini 2.0 Flash 48.15 (0.68) 29.46 (2.03) 51.02 (1.85) 23.67 (0.16) 35.93 (0.42) 42.02 (2.00) 22.35 (0.91)
Open Llama 3.3-70B 34.65 (1.58) 27.80 (4.43) 44.91 (2.31) 37.63 (0.69) 38.70 (1.75) 39.63 (1.51) 45.30 (1.14)
Llama 3.1-8B 34.31 (3.57) 24.56 (0.87) 38.23 (1.70) 30.44 (1.31) 24.61 (1.12) 37.21 (0.95) 32.95 (0.58)
Mixtral-8x7B 38.65 (0.37) 26.82 (0.51) 37.87 (2.36) 31.10 (0.57) 30.35 (2.14) 32.61 (2.42) 26.99 (0.47)
Mistral-7B 28.48 (0.58) 27.71 (0.89) 37.85 (1.01) 34.31 (0.26) 30.01 (1.20) 34.99 (0.63) 32.61 (1.42)
Qwen 3-32B 40.95 (3.35) 30.82 (0.49) 47.55 (1.29) 36.40 (1.36) 36.44 (1.92) 42.87 (0.78) 44.66 (0.42)
Qwen 3-8B 28.18 (1.25) 30.01 (0.58) 40.19 (1.66) 35.55 (2.48) 35.25 (1.01) 36.78 (1.09) 34.70 (0.57)
Qwen 2.5-14B 12.39 (0.28) 21.75 (1.78) 34.40 (2.48) 26.69 (0.45) 24.35 (1.93) 30.57 (3.21) 25.97 (2.29)
Qwen 2.5-7B 12.43 (1.33) 24.14 (0.81) 27.93 (1.16) 22.22 (0.99) 26.82 (0.73) 29.29 (1.85) 23.84 (1.00)
Phi 4-14B 22.86 (1.54) 24.44 (1.44) 33.25 (1.67) 30.10 (0.94) 30.95 (1.66) 28.69 (1.19) 29.50 (0.38)
Phi 4 mini-3.8B 10.47 (0.85) 25.29 (2.08) 31.16 (1.78) 20.99 (0.99) 30.95 (0.63) 24.27 (2.26) 19.75 (0.39)
OLMo 2-7B 24.61 (1.33) 23.37 (0.68) 13.11 (0.87) 24.52 (4.27) 30.35 (0.37) 27.16 (1.12) 28.86 (1.01)
OLMoE-1B-7B 17.97 (0.68) 25.54 (2.18) 6.90 (0.68) 17.41 (6.67) 31.42 (0.28) 23.58 (2.35) 27.71 (0.38)
Tülu 3-8B 24.48 (1.09) 25.59 (0.97) 32.10 (1.17) 30.23 (0.47) 32.35 (0.52) 31.59 (0.69) 29.63 (0.10)
Random Guessing 20.00 20.00 20.00 20.00 20.00 20.00 20.00
@@ 5.3 Results
Overall Performance   Table 3 shows that API-based models outperformed others, likely due to their superior inference capabilities. However, all baseline performances remained below 60%, indicating that current LLMs face substantial limitations when serving as conversational agents in long-term multi-party dialogue settings. Similar trends were observed across the Friends, The Big Bang Theory, and The Office datasets, with detailed results provided in Appendix H.
Extended context windows alone are insufficient for long-term dialogue understanding. As shown in Table 3, models such as GPT-4o-mini, GPT-4.1-nano, and Gemini 2.0 Flash, despite supporting context lengths ranging from 128k to 1M tokens, performed worse than the best retrieval-augmented method, BM25-Session Entire. Only Gemini 2.5 Flash, equipped with both a 1M-token context window and strong reasoning capabilities, achieved the highest overall accuracy of 53.94% under Base LLM setting. These findings suggest that simply increasing the context window is not enough; models must also possess robust reasoning and comprehension capabilities to manage long-term conversations effectively.
Among RAG-based methods, storing the entire session consistently outperforms other history storing methods. As illustrated in Table 3, storing the entire session yields better performance than storing individual utterances or using summarization. This is likely because individual utterances lack sufficient context, and summarization may omit critical information. Moreover, models with strong long-term dialogue understanding (e.g., Gemini 2.5 Flash, Mixtral-8x7B) tended to achieve higher Base-LLM scores, whereas models with strong summarization capabilities (e.g., Llama 3.3-70B) performed best when using the Session Summary method.
Effective memory management is critical for RAG-based agents engaging in long-term multi-party dialogues. We further evaluated model performance in an oracle setting, where agents were
provided with evidence sessions along with timestamps (see Figure 2). As shown in Figure 4, performance in the oracle setting was 10–30% higher compared to that of the best memory management method. This substantial performance gain emphasizes the importance of advanced history storage and retrieval techniques. The complete results of the oracle experiments are provided in Appendix I.
Figure 4: The performance comparison between the oracle setting and the best memory management method.
Table 4: Model performance under different character name settings: (1) using anonymized generic names, (2) using the original character names without anonymization, and (3) swapping the character names (adversarial setting). Performance improves when original names are used and decreases when names are swapped. The full experimental results are provided in Appendix J.
Type Model BM25-Session Entire OpenAI Embedding-Session Entire
Generic Names Original Swapping Names Generic Names Original Swapping Names
API GPT-4o-mini 52.11 55.62 ($\color[rgb]{0,0,1}\uparrow 3.51$}) 45.08 ($\color[rgb]{1,0,0}\downarrow 7.03$}) 45.47 52.81 ($\color[rgb]{0,0,1}\uparrow 7.34$}) 41.38 ($\color[rgb]{1,0,0}\downarrow 4.09$})
Gemini 2.5 Flash 52.92 56.26 ($\color[rgb]{0,0,1}\uparrow 3.34$}) 51.40 ($\color[rgb]{1,0,0}\downarrow 1.52$}) 42.40 50.19 ($\color[rgb]{0,0,1}\uparrow 7.79$}) 43.23 ($\color[rgb]{0,0,1}\uparrow 0.83$})
Gemini 2.0 Flash 51.02 54.34 ($\color[rgb]{0,0,1}\uparrow 3.32$}) 47.81 ($\color[rgb]{1,0,0}\downarrow 3.21$}) 42.02 45.72 ($\color[rgb]{0,0,1}\uparrow 3.70$}) 38.53 ($\color[rgb]{1,0,0}\downarrow 3.49$})
Open Llama3.3-70B 44.91 48.23 ($\color[rgb]{0,0,1}\uparrow 3.32$}) 42.87 ($\color[rgb]{1,0,0}\downarrow 2.04$}) 39.63 43.51 ($\color[rgb]{0,0,1}\uparrow 3.88$}) 38.61 ($\color[rgb]{1,0,0}\downarrow 1.02$})
Mixtral-8x7B 37.87 46.47 ($\color[rgb]{0,0,1}\uparrow 8.60$}) 37.72 ($\color[rgb]{1,0,0}\downarrow 0.15$}) 32.61 41.24 ($\color[rgb]{0,0,1}\uparrow 8.63$}) 31.80 ($\color[rgb]{1,0,0}\downarrow 0.81$})
Qwen 3-32B 47.55 51.72 ($\color[rgb]{0,0,1}\uparrow 4.17$}) 46.32 ($\color[rgb]{1,0,0}\downarrow 1.23$}) 42.87 45.72 ($\color[rgb]{0,0,1}\uparrow 2.85$}) 39.04 ($\color[rgb]{1,0,0}\downarrow 3.83$})
Average 47.73 52.11 ($\color[rgb]{0,0,1}\uparrow 4.38$}) 45.20 ($\color[rgb]{1,0,0}\downarrow 2.53$}) 40.83 46.53 ($\color[rgb]{0,0,1}\uparrow 5.70$}) 38.77 ($\color[rgb]{1,0,0}\downarrow 2.06$})
@@
TKG-based questions are more challenging than fan quiz-style questions, with two-hop reasoning posing particular difficulty. To assess the difficulty levels across different question types, we conducted an error analysis on Gemini-2.5 Flash Base LLM, which showed the highest performance. The results showed that fan quiz-based questions had an accuracy of 62.15%, while TKG-based questions scored lower at 50.83%, highlighting the greater difficulty of TKG-based questions. Breaking down TKG-based questions further, one-hop questions had a performance of 69.19%, whereas two-hop questions had a performance of 19.28%, underscoring the challenge of two-hop questions. Furthermore, even in the oracle setting, while the performance of one-hop questions increased to 83.60%, two-hop questions remained at 51.74%. This indicates that two-hop questions are challenging not only in terms of history retrieval but also in reasoning across the given sessions.
Character anonymization in LongDialQA is essential for fair evaluation of conversational agents. We conducted additional experiments using DialSim on both the original version of LongDialQA (without character anonymization) and the adversarial version where character names were swapped (e.g., Joey $\leftrightarrow$ Monica). As shown in Table 4, performance improved in the original setting, likely because models leveraged pre-trained knowledge alongside dialogue history. These results support the necessity of name anonymization to ensure reliable evaluation. Additionally, performance declined when character names were swapped. This suggests that dialogues with generic names introduce new information, whereas swapped names conflict with pre-trained knowledge, leading to reduced reasoning performance. Detailed results are provided in Appendix J.